Consider a thin spherical shell of radius 14.0 cm with a total charge of 32.0 μC distributed uniformly on its surface. Find the electric field (a) 10.0 cm and (b) 20.0 cm from the center of the charge distribution

 


1
Expert's answer
2021-01-10T18:26:38-0500


Let us draw two imaginary spherical surfaces: inside and outside the shell (see figure). Also let "R = 14cm = 0.14m" be the radius of the shell, "R_1 = 10cm = 0.1m", and "R_2 = 20cm = 0.2m" be the radii of the inner and outer surfaces respectively.

1. Writting down the Gauss's theorem for the inner surface, obtain:


"\oint_{S_1}\mathbf{E_1}d\mathbf{S_1} = \dfrac{Q}{\varepsilon_0}"

where "\mathbf{E_1}" is the electric field on the inner surface "S_1", "\varepsilon_0 = 8.85\times 10^{-12}F/m" is electric constant, and "Q = 0" is the charge enclosed by this surface. Since all charge is located on the shell, "Q = 0", and, thus,


"\mathbf{E_1} = 0"

2. Writting down the Gauss's theorem for the inner surface, obtain:


"\oint_{S_2}\mathbf{E_2}d\mathbf{S_2} = \dfrac{Q}{\varepsilon_0}"

where "\mathbf{E_2}" is the electric field on the outer surface "S_2", and "Q = 32\times 10^{-6}C" is the charge enclosed by this surface. The elecrtic field is constant and perpendicular to the "S_2" at any point (due to spherical symmetry). Thus, the right hand side will be:


"E_2\cdot 4\pi R^2_2 = \dfrac{Q}{\varepsilon_0}"

Expressing "E_2", obtain:


"E_2 = \dfrac{Q}{\varepsilon_0\cdot 4\pi R^2_2}\\
E_2 = \dfrac{32\times 10^{-6}}{8.85\times 10^{-12}\cdot 4\pi \cdot 0.2^2} \approx 7.19\times 10^6N/C"

Answer. a) , b) "7.19\times 10^6N/C".